OpenStudy (anonymous):

what is the answer of the (9+3m+m^2) (3-m) ??

OpenStudy (anonymous):

multiply all terms in the first bracket by 3 and then by -m: =27 + 9m + 3m^2 -9m -3m^2 - m^3 = 27 - m^3

OpenStudy (anonymous):

thank to you very much:D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how about this (4x^2y^2_8xyz+16z^2) ??